What are the various probability distributions, give respective formulas of each type of distribution. Now solve the problem“The probability that at least one of the two independent events occurs is 0.5. Probability thatthe first event occurs but not the second is 3/25. Also the probability that the second event occurs but not the first is 8/25.”Find the probability that none of the two events occurs


1
Expert's answer
2021-06-08T12:56:27-0400

Probability that at least one of A or B occur = "P (A \cup B) = \frac{1}{2}"

Probability that none of the two events occurs "= 1 - P (A \cup B) = 1 - \frac{1}{2}= \frac{1}{2}"